# Sharp Eigenfunction Bounds on the Torus for large $p$

### 11 Mar 2026 v1

@@ 5. Comparison with result with loss
In this section we will catalog the differences between our proofs and the proof in [4].
A more careful integration by parts argument allows us to prove a no loss version of the dispersive type estimate for $G(t,x)$. This is already known by other methods when $q\geq N$. For the $\log$-loss argument had to show the same holds near rationals with very small denominator.
@@
When proving the sharp result, our philosophy is to take any power of $Q$ as long as we do not have an $\epsilon$-loss in the parameters $N$ and $2^{s}$. This gives us worse bounds compared to previous works and prevents us from using more involved calculations. Additionally, our intolerance of losses prevents us from improving over the trivial bound in a few places, most notably (32) and (34). It seems different techniques are needed to improve these bounds.

### 06 Apr 2026 v2

@@ 5. Comparison with result with loss
In this section we will catalog the differences between our proofs and the proofs in [4].
A more careful integration by parts argument allows us to prove a no loss version of the dispersive type estimate for $G(t,x)$. This is already known by other methods when $q\leq N$. For the $\log$-loss argument we had to show the same holds near rationals with very small denominator.
@@
When proving the sharp result, our philosophy is to take any power of $Q$ as long as we do not have an $\epsilon$-loss in the parameters $N$ and $2^{s}$. This gives us worse bounds with respect to $p$ compared to previous works. Additionally, our intolerance of $\epsilon$-losses prevents us from improving over the trivial bound in a few places, most notably (32) and (34). It seems different techniques are needed to improve these bounds.